OpenStudy (anonymous):

Write the equation of a line parallel to the line 2x + 3y = 5 and passes through the point (9, -3).

OpenStudy (anonymous):

slope of parallel lines are same. so, first find the slope of 2x+3y=5 . now you have the slope and a point , apply slope-point form.

OpenStudy (anonymous):

How do I find the slope with only one coodinate?

OpenStudy (anonymous):

you have to find slope of line 2x+3y=5 that will be the slope of required line also. you can find it by converting your eq. in the form y=mx+c then compare.

OpenStudy (anonymous):

3y=-5-2x y= -5/3 - 2/3x compare it with y=mx+c.

OpenStudy (anonymous):

Would the answer be y=-2/3x+5/3?

OpenStudy (anonymous):

thus the equation of the line is (y-(-3))= (-2/3)(x-9).